Debbie traveled by boat 5 miles upstream to fish in her favorite spot. Because of the 4-mph current, it took her 20 minutes longer to get there than to return. How fast will her boat go in still water?
Answer by nerdybill(5411)

Debbie traveled by boat 5 miles upstream to fish in her favorite spot. Because of the 4-mph current, it took her 20 minutes longer to get there than to return. How fast will her boat go in still water?
.
Let t = time going upstream (with current)
and b = speed of boat in still waters
.
from: "Debbie traveled by boat 5 miles upstream to fish in her favorite spot."
5 = t(b+4)
from: "Because of the 4-mph current, it took her 20 minutes longer to get there than to return."
5 = (t + 1/3)(b-4)
.
5 = t(b+4)
t = 5/(b+4)
.
Substitute the above into equation 2:
5 = (t + 1/3)(b-4)
5 = (5/(b+4) + 1/3)(b-4)
15(b+4) = (15 + (b+4))(b-4)
15(b+4) = (b+19)(b-4)
15b+60 = b^2-4b+19b-76
15b+60 = b^2+15b-76
60 = b^2-76
136 = b^2
11.662 mph = b (speed of boat in still water)